An object of mass moves round a circle of radius 6 m if the if the period of motion is π calculate the force towards the centre

1
Expert's answer
2021-11-14T17:09:23-0500

Assume that "m=5\ kg" . So, we have


"F=mv^2/r=m\omega^2r=m(\frac{2\pi}{T})^2r=m(\frac{2\pi}{\pi})^2r="


"=4mr=4\cdot5\cdot6=120\ (N)" . Answer